OpenStudy (luigi0210):

This is probably easy for some of ya but I just don't understand word problems.. so help please? :) Free Undamped Motion

OpenStudy (luigi0210):

A mass weighing 4 pounds is attached to a spring whose spring constant is 16 lb/ft. What is the the period of simple harmonic motion?

OpenStudy (yanasidlinskiy):

Isn't there supposed to be some kind of formula?

OpenStudy (yanasidlinskiy):

I know one just not sure about it.

OpenStudy (anonymous):

\[F = k \times x\]

OpenStudy (anonymous):

Force = Spring Constant times Displacement, is that right? I am also not sure. :P

OpenStudy (luigi0210):

That's part of the formula

OpenStudy (yanasidlinskiy):

Not sure. There's a different one I know. Just a sec. Let me type it out. Gawd. My brains will be fried by the end of the dayXD

OpenStudy (yanasidlinskiy):

\[\frac{ d^2x }{ dt^2 } + \omega^2x = F _{0}\] This may seem confusing.....

OpenStudy (luigi0210):

Yes, that one ^ I'm just having trouble plugging in the information..

OpenStudy (yanasidlinskiy):

Yay! I got it!! I'm so proud of myselfXD Anywho..Lets see where everything goesXD

OpenStudy (anonymous):

Just doing this one for myself only, don't look at it. :P \[X(s)[s^2 + \omega^2] = F(s) \\ \frac{X(s)}{F(s)} = \frac{1}{s^2 + \omega^2}\]

OpenStudy (anonymous):

My phone does not have this option.. :P

OpenStudy (yanasidlinskiy):

LOL. Quit spamming this question. I'm thinking........@e.mccormick shall know:P

OpenStudy (yanasidlinskiy):

That didn't work well. @e.mccormick

OpenStudy (anonymous):

Is not the equation supposed to be: \[M \frac{d^2x}{dt^2} + kx = F\]

OpenStudy (anonymous):

Nice try @YanaSidlinskiy

OpenStudy (luigi0210):

Yea, but we divided by m to make it \(\Large x''+\omega ^2 =0 \) Where omega is just a sub for k/m

OpenStudy (luigi0210):

*x

OpenStudy (anonymous):

I think then omega is sub for root(k/m) and not simply k/m ?

OpenStudy (anonymous):

Then it should be simple I guess..

OpenStudy (yanasidlinskiy):

Eta, there's a thing called where "we all make mistakes, So *******:) " Stop spamming the post, Eta ;) Lets get to work. This is gonna be fun.

OpenStudy (luigi0210):

My textbook gives me \(\Large \frac{d^2x}{dt^2}+\omega ^2 x=0 \) So I'm just following the book :P

OpenStudy (anonymous):

\[\omega = \sqrt{\frac{k}{m}} = \sqrt{\frac{16}{4}}\]

OpenStudy (anonymous):

Why are you so serious yana? If question is a fun, then you can solve it by having fun too. :)

OpenStudy (anonymous):

You just need to adjust units now..

OpenStudy (yanasidlinskiy):

Oh, what else does the book tell you? I'm sure there's some answers in the back of the book. Also try and look at the practice problems if it gives you any. I'm sure it does. Eta, it's sarcasm ;)

OpenStudy (anonymous):

Yeah, I know, never to argue with girls, at the end, you are gonna lose..!!

OpenStudy (luigi0210):

the period they give is \(\Large \frac{ \sqrt{2} \pi}{8} \)

OpenStudy (yanasidlinskiy):

So wait. Which formula are we using? And Luigi, so what does 2 and 8 stand for? My brains are fried already, LOL! Eta,

OpenStudy (anonymous):

Sorry, but what is relation between lbs and pounds?

OpenStudy (luigi0210):

You guys confuse me more -.- Let's see if @hartnn is alive to help :3

OpenStudy (luigi0210):

lb is short for pounds

OpenStudy (anonymous):

Oh..!!

OpenStudy (yanasidlinskiy):

Ya, It's for pounds. Like how much you weigh;) Hah! @hartnn ova here.

hartnn (hartnn):

period is 2pi/omega

hartnn (hartnn):

i see we don't get omega as 8 sqrt 2 hmm

OpenStudy (anonymous):

How come units of Spring constant is pounds/feet?

OpenStudy (anonymous):

I think, it must be Newton/feet or Newton/Meter in SI..

OpenStudy (luigi0210):

Cause this is America xD

hartnn (hartnn):

k = mg/L , since we don't have L, this is not useful...

OpenStudy (yanasidlinskiy):

Maybe @ganeshie8 might know?

OpenStudy (luigi0210):

they told me to use \(32 ft/s^2 \) for something but I don't remember what

hartnn (hartnn):

that is for 'g'

OpenStudy (anonymous):

That is g I think..

OpenStudy (anonymous):

g has SI units of m/sec^2

OpenStudy (anonymous):

Oh, I think now I got it.. :P

hartnn (hartnn):

"weighing 4 ponds"!

hartnn (hartnn):

so thats mg mg =4

hartnn (hartnn):

find m from here

hartnn (hartnn):

use m = 4/32 = 1/8 and try again?

hartnn (hartnn):

you'll get it :)

OpenStudy (anonymous):

mg = 4? are you sure?

hartnn (hartnn):

yes :) weight is mg mass is m

OpenStudy (anonymous):

:)

hartnn (hartnn):

I am sure Lui can do all the substitutions and get his correct answer :)

OpenStudy (anonymous):

Finally, got it.. :P

OpenStudy (luigi0210):

It took me awhile but I finally figured out what you meant, thanks guys for all your time and help :)

hartnn (hartnn):

welcome ^_^

OpenStudy (yanasidlinskiy):

:) Yw.XD

OpenStudy (anonymous):

\(\dagger\)